# Omni-Perception: Omnidirectional Collision Avoidance for Legged Locomotion in Dynamic Environments

### v1

@@ Abstract
Agile locomotion in complex 3D environments requires robust spatial awareness to safely avoid diverse obstacles such as aerial clutter, uneven terrain, and dynamic agents. Depth-based perception approaches often struggle with sensor noise, lighting variability, computational overhead from intermediate representations (e.g., elevation maps), and difficulties with non-planar obstacles, limiting performance in unstructured environments. In contrast, direct integration of LiDAR sensing into end-to-end learning for legged locomotion remains underexplored. We propose Omni-Perception, an end-to-end locomotion policy that achieves 3D spatial awareness and omnidirectional collision avoidance by directly processing raw LiDAR point clouds. At its core is PD-RiskNet (Proximal-Distal Risk-Aware Hierarchical Network), a novel perception module that interprets spatio-temporal LiDAR data for environmental risk assessment. To facilitate efficient policy learning, we develop a high-fidelity LiDAR simulation toolkit with realistic noise modeling and fast raycasting, compatible with platforms such as Isaac Gym, Genesis, and MuJoCo, enabling scalable training and effective sim-to-real transfer. Learning reactive control policies directly from raw LiDAR data enables the robot to navigate complex environments with static and dynamic obstacles more robustly than approaches relying on intermediate maps or limited sensing. We validate Omni-Perception through real-world experiments and extensive simulation, demonstrating strong omnidirectional avoidance capabilities and superior locomotion performance in highly dynamic environments. We will open-source our code and models.
@@
Keywords: Legged Robot, Locomotion, Reinforcement Learning, Collision Avoidance, LiDAR Perception
@@ 1 Introduction
Meanwhile, LiDAR sensors have become instrumental in fields such as autonomous driving, offering dense, lighting-invariant 3D measurements, and manipulation [19, 20, 21, 22, 23, 24, 25]. LiDAR delivers rich geometric information ideal for navigating spatially complex environments. Despite its promise, the integration of LiDAR—especially direct point cloud processing—into end-to-end learning for legged locomotion remains underexplored [26, 27, 28]. This gap arises from the challenges of real-time point cloud processing within fast control loops, and from difficulties in accurately modeling LiDAR physics for effective sim-to-real transfer [29].
@@
    Novel LiDAR Perception Network (PD-RiskNet): We introduce a hierarchical network architecture specifically designed to process spatio-temporal LiDAR point clouds, differentiating between proximal and distal regions to effectively quantify environmental risks for locomotion.
@@ 4.1 LiDAR fidelity Evaluation
As illustrated in Figure. 3, we validate our LiDAR simulator’s fidelity. We compare its output to real data from a Unitree G1 robot using a Livox Mid-360 sensor and Isaac Sim LiDAR using the same horizontal angle and vertical angle. Our simulation includes the sensor’s non-repetitive scan pattern and robot self-occlusion. The simulated and real point distributions and structures appeared similar. This result indicates high simulator fidelity, which helps reduce the sim2real perception gap for LiDAR policies.You can find more lidar model scan patterns in the Appendix.F
@@ Acknowledgments
If a paper is accepted, the final camera-ready version will (and probably should) include acknowledgments. All acknowledgments go at the end of the paper, including thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
@@ A.1 Rewards
Velocity Tracking with Avoidance ($r_{vel,avoid}$) $\exp\{{-\beta_{va}||\mathbf{v}_{t}-(\mathbf{v}_{t}^{\text{cmd}}+\mathbf{V}_{\text{avoid},t})||^{2}}\}$ 2
@@
[…]
link collision $||Force^{\text{PenltyLink}}_{xy}||^{2}$ $-0.02$
joint limit violation $\mathbf{1}_{q_{i}>q_{\max}||q_{i}<q_{\min}}$ $-0.2$
[…]
@@ Appendix D Domain Randomization Details
Added Mass (kg) $-1.0$ $5.0$
Payload Mass (kg) $-1.0$ $3.0$
Center of Mass x (m) $-0.1$ $0.1$
Center of Mass y (m) $-0.15$ $0.15$
Center of Mass z (m) $-0.2$ $0.2$
@@
Joint Calibration Offset (rad) $-0.02$ $0.02$
Gravity Offset (m/s2) $-1.0$ $1.0$
@@ Appendix E Taichi Lidar Efficiency
Appendix F Lidar Pattern
(a) Mid360 Pattern
(b) Mid40 Pattern
(c) Mid70 Pattern
(d) Avia Pattern
(e) Tele Pattern
(f) HDL64 Pattern
(g) VLP32 Pattern
(h) OS128 Pattern
Figure 7: Various LiDAR Scan Patterns

### v2

@@ Abstract
Agile locomotion in complex 3D environments requires robust spatial awareness to safely avoid diverse obstacles such as aerial clutter, uneven terrain, and dynamic agents. Depth-based perception approaches often struggle with sensor noise, lighting variability, computational overhead from intermediate representations. In contrast, direct integration of LiDAR sensing into end-to-end learning for legged locomotion remains underexplored. We propose Omni-Perception, an end-to-end locomotion policy that achieves 3D spatial awareness and omnidirectional collision avoidance by directly processing raw LiDAR point clouds. At its core is PD-RiskNet (Proximal-Distal Risk-Aware Hierarchical Network), a novel perception module that interprets spatio-temporal LiDAR data for environmental risk assessment. To facilitate efficient policy learning, we develop a high-fidelity LiDAR simulation toolkit with realistic noise modeling and fast raycasting, compatible with platforms such as Isaac Gym, Genesis, and MuJoCo, enabling scalable training and effective sim-to-real transfer. Learning reactive control policies directly from raw LiDAR data enables the robot to navigate complex environments with static and dynamic obstacles more robustly than approaches relying on intermediate maps or limited sensing. We validate Omni-Perception through real-world experiments and extensive simulation, demonstrating strong omnidirectional avoidance capabilities and superior locomotion performance in highly dynamic environments.
@@
Keywords: Legged Robot, Locomotion, Reinforcement Learning, LiDAR Perception
@@ 1 Introduction
Meanwhile, LiDAR sensors have become instrumental in fields such as autonomous driving, offering dense, lighting-invariant 3D measurements, and manipulation [19, 20, 21, 22, 23, 24, 25, 26]. LiDAR delivers rich geometric information ideal for navigating spatially complex environments. Despite its promise, the integration of LiDAR—especially direct point cloud processing—into end-to-end learning for legged locomotion remains underexplored [27, 28, 29]. This gap arises from the challenges of real-time point cloud processing within fast control loops, and from difficulties in accurately modeling LiDAR physics for effective sim-to-real transfer [30].
@@
    Novel LiDAR Perception Network (PD-RiskNet):
@@ 4.1 LiDAR fidelity Evaluation
As illustrated in Figure. 3, we validate our LiDAR simulator’s fidelity. We compare its output to real data from a Unitree G1 robot using a Livox Mid-360 sensor and Isaac Sim LiDAR using the same horizontal angle and vertical angle. Our simulation includes the sensor’s non-repetitive scan pattern and robot self-occlusion. The simulated and real point distributions and structures appeared similar. This result indicates high simulator fidelity, which helps reduce the sim2real perception gap for LiDAR policies.You can find more lidar model scan patterns in the Appendix.LABEL:appendix:Lidar_pattern
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62306257) and the Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2023A03J0008), Education Bureau of Guangzhou Municipality.
@@ A.1 Rewards
[…]
@@
z velocity $v_{z}^{2}$ $3\text{\times}{10}^{-4}$
foot stumble $Force^{\text{foot}}_{xy}||^{2}$ $2\text{\times}{10}^{-2}$
link collision $Force^{\text{PenltyLink}}_{xy}||^{2}$ 0.02-0.02
joint limit violation $\mathbf{1}_{q_{i}>q_{\max}||q_{i}<q_{\min}}$ 0.2-0.2
joint torques $\boldsymbol{\tau}||^{2}$ $1\text{\times}{10}^{-6}$
joint velocities $\dot{\mathbf{q}}||^{2}$ $1\text{\times}{10}^{-6}$
joint accelerations $\ddot{\mathbf{q}}||^{2}$ $2.5\text{\times}{10}^{-7}$
action smoothing $\mathbf{a}_{t-1}-\mathbf{a}_{t}||^{2}$ $5\text{\times}{10}^{-3}$
action smoothing rate $\mathbf{a}_{t-2}-2\mathbf{a}_{t-1}+\mathbf{a}_{t}||^{2}$ $5\text{\times}{10}^{-3}$
@@ Appendix D Domain Randomization Details
Added Mass (kg) 1.0-1.0 $5.0$
Payload Mass (kg) 1.0-1.0 $3.0$
Center of Mass x (m) 0.1-0.1 $0.1$
Center of Mass y (m) 0.15-0.15 $0.15$
Center of Mass z (m) 0.2-0.2 $0.2$
@@
Joint Calibration Offset (rad) 0.02-0.02 $0.02$
Gravity Offset (m/s2) 1.0-1.0 $1.0$